what are the two binomial factors of 6s^2+40s-64
Answer by stanbon(26297)

what are the two binomial factors of 6s^2+40s-64
--------
= 2(3s^2+20s-32)
---
Think of two numbers whose product is -32*3 = -96
and whose sum is 20.
The numbers are 24 and -4
Rewrite the problem:
= 2(3s^2+24s-4s-32)
Factor the 1st two and the last two terms of
the quadratic separately.
=2(3s^2+24s-4s-32)
=2(3s(s+8)-4(s+8))
Factor again to get:
---
= 2(3s-4)(s+8)
====================
Cheers,
Stan H.
===================